# Supervised-Contrastive Loss Learns Orthogonal Frames and Batching Matters

### v1

@@ 3 SCL learns OF geometries: Empirical findings
Experimental setup. We study the behavior of models trained with SCL under, 1) $R$-STEP imbalance having $k/2$ majority classes with $n_{\text{maj}}$ datapoints per class and $k/2$ minority classes with $n_{\min}=\frac{n_{{\text{maj}}}}{R}$ datapoints per class, and 2) $R$-Long-tailed (LT) imbalance where the number of training datapoints exponentially decreases across classes such that $n_{c}=n_{1}R^{-(c-1/k-1)}$, for $c\in[k]$. Regardless of the imbalance ratio $R$, we ensure that $n_{c}\geq 2$ by adding a vertically flipped version of each image as a method of batch duplication. Unless explicitly stated, we do not perform any additional data augmentation on the datasets, and use a batch size of $1024$ with random reshuffling. We minimize SCL using Stochastic Gradient Descent (SGD) with a constant learning rate of $0.1$, and momentum set to $0.9$ with no weight decay. Furthermore, as in [Gra+21, Kho+20], we set the temperature parameter $\tau=0.1$. We empirically find that convergence to OF is not highly dependent on $\tau$ for values near $0.1$, yet the specific choice affects the speed of convergence. Besides this, [Kho+20, ] have found that $\tau=0.1$ leads to the best generalization performance. In all cases, network architectures (ResNet-18 etc.) normalize features to unit norm before the output as discussed in [Kho+20]. Link to the codebase to reproduce experiments is provided here.
@@
Metrics. Since in all experiments we observe the within-class variations of the last-layer embeddings ($\mathbf{h}_{i}$) become negligible towards the end of training (see NC plots in Fig. 10 in the appendix), we focus here on the geometry of class-mean embeddings $\mathbf{M}$. Thus, to measure the distance of learned embedding to the OF geometry, we compute the distance metric $\Delta_{\mathbf{G}_{\mathbf{M}}}:=\|\frac{\mathbf{G}_{\mathbf{M}}}{\|\mathbf{G}_{\mathbf{M}}\|_{F}}-\frac{\mathds{I}_{k}}{\|\mathds{I}_{k}\|_{F}}\|_{F}$. In addition, we measure the pairwise cosine similarity $\alpha_{\text{sim}}(c,c^{\prime}):=\nicefrac{{\bm{\mu}_{c}^{\top}\bm{\mu}_{c^{\prime}}}}{{\|\bm{\mu}_{c}\|\|\bm{\mu}_{c^{\prime}}\|}}$ for $c\neq c^{\prime}\in[k]$. Finally, for ease of comparison with results of [Gra+21, ], we also compute the corresponding angular distance $\Delta_{\text{ang}}:=1-\nicefrac{{1}}{{\pi}}\arccos(\alpha_{\text{sim}}(c,c^{\prime}))$.
[…]
Figure 3: $\mathbf{G}_{\mathbf{M}}$ at last epoch (350) for ResNet-18 trained on $R$-STEP imbalanced MNIST. See text for details.
@@
Observations. In Fig. 2, we plot the distance $\Delta_{\mathbf{G}_{\mathbf{M}}}$ between the learned feature-embeddings and the OF geometry as training progresses. Across our deep-net experiments, we consistently observe that the learned features during training converge to the OF geometry, irrespective of the imbalance level $R$ of the training set and the imbalance pattern (STEP or LT). This suggests that the feature geometry learned by SCL is invariant to the training label distribution.
@@
This property contrasts with that of CE and MSE loss which are known to be sensitive to imbalances [Thr+22, Fan+21, Liu+23, Beh+23, Dan+23]. Additionally, for different values of $R$, there is no significant difference between the speed of convergence, unlike the CE loss [Thr+22], where it is empirically observed that the rate worsens with larger imbalance levels.
Lastly, Fig. 3 compares the heatmap of $\mathbf{G}_{\mathbf{M}}$ at last epoch 350 learned by SCL and CE losses over $R$-STEP imbalanced MNIST. For better comparison, in each experiment, we normalize $\mathbf{G}_{\mathbf{M}}$ by the maximum entry (i.e., we normalize the class-mean features to have a maximum norm of $1$). We observe that in case of CE with high imbalance ratio $R$, the class-mean embeddings representing different minority classes are closer to each other (have smaller angles) compared to that of SCL; see Fig. 3, $R=100$.
Remark 1.
Contrary to previous claims in the literature, our findings challenge the notion that the embeddings geometry of SCL becomes asymmetric under imbalances, as suggested by [Zhu+22, ]. Instead, our results indicate that the geometry remains an OF, regardless of the presence of class imbalance. It is important to highlight that the primary focus and main contribution of [Zhu+22] revolve around their proposed variant of SCL, called the balanced contrastive loss (BCL), which they empirically demonstrate achieves state-of-the-art performance in long-tailed image classification tasks. Nevertheless, our findings raise an intriguing research question that challenges the justification for BCL given in [Zhu+22], particularly the claim that it enforces symmetric geometric structures unlike SCL. In particular, this prompts further investigation into understanding the underlying factors that make BCL more resilient to distributional variations compared to SCL.
@@ 4 Theoretical justification: SCL with non-negativity constraints
In this section, we analytically justify the convergence of the embeddings learned by SCL to the OF geometry. For our theoretical analysis, we use the Unconstrained Features Model (UFM) [MPP20, Fan+21, Gra+21, Ji+21, LS20, TB22], where we treat the last-layer features ${\mathbf{H}}$ as free variables, removing the dependence to the network parameters ${\bm{\theta}}$. However, we refine the UFM, and consider the SCL minimization over the non-negative orthant to accommodate for the presence of ReLU activations in the last layer of many common architectures.444[TB22] has also studied incorporating the ReLU activation in the UFM, albeit focusing on MSE loss. We further constrain the embeddings $\mathbf{h}_{i}$ to be normalized as we saw this is inherent in the SCL optimization [Kho+20]. We prove, irrespective of the labels distribution, that the global optimizers form an OF. Formally, consider the following refined UFM for SCL, which we call $\text{UFM}_{+}$ for convenience:
@@
Recall ${\mathbf{H}}\geq 0$ denotes entry-wise non-negativity. We begin our analysis by considering the full-batch version of the loss in Sec. 4.1, and subsequently extend our results to the SCL minimized on mini-batches as in (3) in Sec. 4.2. We provide specific conditions that mini-batches must satisfy in order to have the same global optimum as the full-batch version.
@@ 4.1 Full-batch SCL
Remark 2.
@@
Figure 4: Full-batch SCL converges to the lower bound (dashed lines) computed in Thm. 1 during training. See text for details.
We defer the details of the proof to Sec. A in the appendix. The bound relies on successive uses of Jensen’s inequality and the fact that for feasible ${\mathbf{H}}$, each pair of training samples $i,j\in[n]$ satisfies $0\leq\mathbf{h}_{i}^{\top}\mathbf{h}_{j}\leq 1$. We complete the proof by verifying that equality is attained only if $\mathbf{h}_{i}^{\top}\mathbf{h}_{j}=1$ when $y_{i}=y_{j}$ and $\mathbf{h}_{i}^{\top}\mathbf{h}_{j}=0$ when $y_{i}\neq y_{j}$. In other words, to achieve the optimal cost, the features with similar labels must align (NC) and the class-mean features should form an OF.
4.2 Mini-batch SCL
Thm. 1 shows that minimizing SCL over the training set as a single batch recovers an OF in the feature space, regardless of the training label distribution. However, in practice, SCL optimization is performed over batches chosen from training set [Kho+20, Gra+21]. Specifically, we have a set of batches $\mathcal{B}$ and we compute the loss on each batch $B\in\mathcal{B}$ as in (3). While in the full-batch SCL (4), all pairs of training samples interact with each other, in the mini-batch version, two samples $(i,j)$ interact only if there exists a batch $B\in\mathcal{B}$ such that $i,j\in B$. The natural question is whether the mini-batch construction impacts the embeddings learned by SCL. In this section, we study the role of batching on the embeddings geometry more closely.
where recall $n_{B,c}=|\{i:i\in B,y_{i}=c\}|$. The following theorem is an extension of Thm. 1 for the mini-batch SCL (10). The proof proceeds similarly to that of Thm. 1, where the loss over each batch can be independently bounded from below. Interestingly, due to the orthogonality of the optimal embeddings in every batch, the lower bound for each batch can be achieved by the common minimizer. The detailed proof is deferred to Sec. B in the appendix.
Theorem 2 (Mini-batch SCL minimizers).
@@ 4.2 Mini-batch SCL
Moreover, equality holds if and only if for every $B\in\mathcal{B}$ and every pair of samples $i,j\in B$, $\mathbf{h}_{i}^{\top}\mathbf{h}_{j}=0$ if $y_{i}\neq y_{j}$ and $\mathbf{h}_{i}=\mathbf{h}_{j}$ if $y_{i}=y_{j}$.
Figure 5: A non-OF optimal feature geometry minimizing the mini-batch SCL (10) with =$\mathcal{B}=\{B_{1},B_{2}\}$. Samples with different labels are marked with different colors. See text for details.
Global minimizer geometry may not be unique. It is easy to verify that any ${\mathbf{H}}$ following the OF geometry achieves the lower-bound and it is also a global optimizer of the mini-batch SCL. However, depending on the choice of $\mathcal{B}$, the lower bound can possibly be attained by other optimal embeddings that do not necessarily satisfy NC or orthogonality. Hence the global optimizer may not have a unique geometry.
@@ 5 Batching matters
By the fact that the OF geometry may not be the unique solution for the $\text{UFM}_{+}$ with mini-batch SCL (10), here we study the role of mini-batches for optimizing SCL to avoid ambiguous solutions with different geometries. In Sec. 5.1, we first identify necessary and sufficient conditions for a batching strategy to yield a unique solution geometry to the global minimizers. Then, in Sec. 5.2, using this result, we propose a simple yet effective scheme that provably succeeds in improving the convergence to an OF.
@@ 5.1 When is OF geometry the unique minimizer?
With this definition, we can now state the necessary and sufficient conditions on $\mathcal{B}$ for the minimizer of the mini-batch SCL to be uniquely OF.
@@
Corollary 2.1 serves as a check to ascertain whether a given batching scheme yields a unique global minimizer geometry or not. It also provides guidance for designing mini-batches. We elaborate on this aspect in the following section.
5.2 Batch-binding improves convergence
Corollary 2.1 shows that an arbitrary set of mini-batches is not guaranteed to have OF geometry as its minimizing configuration in the embedding space. To enforce the OF geometry as the unique minimizer, we introduce a simple scheme with low computational overhead as follows.
@@ 5.2 Batch-binding improves convergence
Algorithm 1 increases the size of every mini-batch by a constant value $k$. This value is typically smaller than the batch sizes used to train SCL which could range from $2048$ to $6144$ [Kho+20]. While adding small computational overhead, this method guarantees that $\mathcal{B}^{\prime}$ satisfies the conditions of Cor. 2.1 and SCL learns an OF geometry. We conduct a series of experiments to illustrate the impact of batch binding on convergence to OF.
Geometry convergence for different batching schemes.  To verify the role of batch selection on the learned features, we train ResNet-18 on CIFAR-10 under three batching scenarios: 1) No batch-shuffling: a mini-batch set $\mathcal{B}$ that partitions the training examples once at initialization and is held constant across training epochs. 2) Batch-shuffling: the examples are divided into mini-batch partition, with a random reshuffle of the examples at every epoch. 3) No batch-shuffling + batch-binding: we construct a fixed partition of examples into mini-batches. Further, to each mini-batch we append a fixed set of $k$ binding examples. This remains constant throughout the training.
@@
Fig. 6 shows the distance of learned embeddings to the OF geometry for the three batching scenarios mentioned above. In case of the fixed partition batching (left), the features do not converge to OF, especially with large imbalance ratios. This behavior is consistent with the conditions identified by Thm. 2, because in a typical random partition, the corresponding induced subgraphs are not necessarily connected. On the other hand, when the batching is performed with a randomly reshuffled data ordering (center), the geometry converges more closely to the OF, albeit with certain epochs deviating significantly from the convergence direction. We hypothesize that random reshuffling creates an opportunity for different examples to interact sufficiently, thus, when trained long enough, to converge close to an OF geometry.
Finally, with the third batching strategy (right), by adding the binding examples to the batches, even without shuffling the samples at every epoch, we observe a consistently fast convergence to the global optimizer, OF. These observations provide strong evidence supporting our analysis of batching predicting the behavior of SCL trained deep-nets. Moreover, in Table 1 in the appendix, we present experimental evidence in suggesting that batch binding does not negatively effect the balanced test accuracy of Nearest Class Center (NCC) classifier. This observation, paired with the impact of binding examples on OF convergence, motivates further analysis of the relationship between OF geometry and test accuracy.
[…]
Improving convergence for less powerful models.  When conducting our experiments for Fig. 1, we have noticed that DenseNet-40 converges slower compared to ResNet-18. A reason for this may be related to the very different complexities of the two models: ResNet-18 has substantially more trainable parameters compared to DenseNet-40. In an attempt to improve the convergence speed, we reduce the batch size for DenseNet experiments to $128$ to increase the number of SGD iterations. We also train DenseNet for $500$ epochs instead of $350$ while reducing the learning rate from $0.1$ to $0.01$ at epoch $300$. Yet, we observe that with much smaller batch sizes, the embedding geometry does not always converge to the OF geometry, especially when training with high imbalance ratios. Specifically when training with randomly reshuffled batches, there is a higher chance that the examples do not interact with each other even if trained for long, suggesting that the optimal solution for all batches is not necessarily OF. We hypothesize that this likelihood increases when using smaller batch sizes during training. In order to combat this effect, we ran the DenseNet experiments with the addition of binding examples to every batch. As shown in Fig. 7, we find that adding such binding examples significantly improves the convergence to OF, particularly when training on more complex datasets (CIFAR10 compared to MNIST) and under more severe imbalances (STEP imbalance with $R=50,100$). These results emphasize the impact of batch-binding and provide further evidence regarding our claims in Sec. 4.2. While the convergence values of $\Delta_{\mathbf{G}_{\mathbf{M}}}$ are higher when compared to ResNet-18, we deem them reasonable considering the difference in the number of parameters between the two architectures. Similar to ResNet-18, we provide geometric convergence results for DenseNet-40 on different datasets in Fig. 19 in the appendix.
[…]
Improving convergence for more complex datasets.  In addition to the benchmark datasets previously explored, we consider the convergence of ResNet-18 embeddings to the OF geometry when trained with CIFAR100. The models are trained for $500$ epochs with a constant learning rate of $0.1$ and a batch size of 1024. With a large number of classes ($k=100$ for CIFAR100), it becomes increasingly more likely for the randomly reshuffled batches to not allow sufficient interactions between examples. Particularly for large imbalance ratios (e.g., $R=100,50$), since the number of samples in each minority class could become as low as $5$-$10$, some batches might never encounter examples from minority classes. However, including the binding examples in each batch can improve the convergence of the feature geometry to OF. In Fig. 8 we measure the convergence of ResNet-18 trained both with and without the batch-binding. Batch-binding considerably expedites and improves the convergence to the OF. This is particularly noticeable for STEP-Imbalance and large imbalance ratios, where half of the classes are minorities, thus the inclusion of binding examples becomes crucial in ensuring that our theoretical requirements for reaching the OF are satisfied.
On the convergence of batch-shuffling to OF.  It can be readily observed that a fixed mini-batch partition without shuffling (No batch-shuffling in the experiments above) does not satisfy the conditions of Cor. 2.1. Consequently, OF is not the unique minimizer geometry in such scenarios. This is clearly manifested by the inadequate convergence levels observed in our experiments, as depicted for example in Fig. 6. In contrast, in our experiments, when the examples are randomly shuffled prior to partitioning into mini-batches at each epoch (which we call Batch Shuffling), the convergence behavior to OF shows a significant improvement compared to the case of No batch-shuffling. The observed improvement can be attributed to the fact that shuffling enables interactions among examples across epochs. To elaborate on this notion, we present a formalization of batch shuffling below.
Remark 3.
6 Discussion
6.1 More related work

### v2

@@ 3 SCL with ReLU learns OF geometries: Empirical findings
We study the behavior of models trained with SCL under, 1) $R$-STEP imbalance having $k/2$ majority classes with $n_{\text{maj}}$ examples per class and $k/2$ minority classes with $n_{\min}=\nicefrac{{n_{{\text{maj}}}}}{{R}}$ examples per class, and 2) $R$-Long-tailed (LT) imbalance where the number of training datapoints exponentially decreases across classes such that $n_{c}=n_{1}R^{-(c-1/k-1)}$, for $c\in[k]$. Regardless of the imbalance ratio $R$, we ensure that $n_{c}\geq 2$ by adding a vertically flipped version of each image as a method of batch duplication. Unless stated, we do not perform any additional data augmentation on the datasets, and use a batch size of $1024$ with random reshuffling.
Finally, when studying the impact of ReLU on generalization, we use a Nearest Center Classifier (NCC) on the output features to evaluate model performance. For such experiments, following the setup of [Kho+20], we employ a projection head by adding a 2 layer non-linear MLP (with or without ReLU at the last layer) and compare the test accuracy under difference imbalance ratios.
Imbalance Ratio (R) w/o ReLU w/ ReLU
1 72.17 $\pm$ 0.23 72.32 $\pm$ 0.60
10 (Step) 56.58 $\pm$ 0.50 58.16 $\pm$ 0.76
10 (LT) 57.10 $\pm$ 1.04 57.88 $\pm$ 1.05
100 (Step) 43.49 $\pm$ 0.30 43.80 $\pm$ 0.25
100 (LT) 37.19 $\pm$ 2.50 39.71 $\pm$ 0.09
Table 1: Test accuracy comparison for a ResNet-18 trained using SCL on CIFAR100 with and without ReLU after the projection head. We use NCC for classification. Note that the addition of ReLU does not compromise the accuracy. For details, see Sec. E.7 in the appendix.
@@
Metrics. Since in all experiments we observe the within-class variations of the last-layer embeddings ($\mathbf{h}_{i}$) become negligible towards the end of training (see NC plots in Fig. 8 in the appendix), we focus here on the geometry of class-mean embeddings $\mathbf{M}$. Thus, to measure the distance of learned embedding to the OF geometry, we compute the distance metric $\Delta_{\mathbf{G}_{\mathbf{M}}}:=\|\frac{\mathbf{G}_{\mathbf{M}}}{\|\mathbf{G}_{\mathbf{M}}\|_{F}}-\frac{\mathds{I}_{k}}{\|\mathds{I}_{k}\|_{F}}\|_{F}$.
@@
Observations on Geometry. In Fig. 2, we plot the distance $\Delta_{\mathbf{G}_{\mathbf{M}}}$ between the learned feature-embeddings and the OF geometry as training progresses. We consistently observe that the learned features during training converge to the OF geometry, irrespective of the imbalance level $R$ of the training set and the imbalance pattern (STEP or LT). This suggests that the feature geometry learned by SCL is invariant to the training label distribution.
@@
Interestingly, the invariance that is revealed by our study is distinct for SCL with ReLU as opposed to the case of SCL without ReLU, and that of CE and MSE loss. CE and MSE are known to be sensitive to imbalances [Thr+22, Fan+21, Liu+23, Beh+23, Dan+23]. Additionally, for different values of $R$, there is no significant difference between the speed of convergence, unlike the CE loss [Thr+22], where it is empirically observed that the rate worsens with larger imbalance.
Generalization. To ensure that ReLU does not yield any adverse effects on test accuracy, we trained ResNet-18 (with a projection head) with and without ReLU on CIFAR10, CIFAR100, Tiny Imagenet and evaluated the balanced test accuracy when training under label imbalance. Following [Zhu+22, Kho+20, Che+20], we consider the features before a projection head for evaluating the test accuracy of the models. Our results, e.g. in Tab. 1, indicate that the addition of ReLU does not compromise test accuracy. For a detailed discussion on the results see Sec. E.7 in the appendix.
@@ 4 Theoretical justification: SCL with non-negativity constraints
In this section, we analytically justify the convergence of the embeddings learned by SCL to the OF geometry. For our theoretical analysis, we use the Unconstrained Features Model (UFM) [MPP20, Fan+21, Gra+21, Ji+21, LS20, TB22], where we treat the last-layer features ${\mathbf{H}}$ as free variables, removing the dependence to the network parameters ${\bm{\theta}}$. However, we refine the UFM, and consider the SCL minimization over the non-negative orthant to accommodate for the presence of ReLU activations in the last layer.555[TB22] (and several follow ups, e.g. [SML23]) has also studied incorporating the ReLU activation in the UFM, albeit focusing on MSE loss. Following [Kho+20, ], we further constrain the embeddings $\mathbf{h}_{i}$ to be normalized. We prove, irrespective of the labels distribution, that the global optimizers form an OF. Formally, consider the following refined UFM for SCL, which we call $\text{UFM}_{+}$ for convenience:
@@
Recall ${\mathbf{H}}\geq 0$ denotes entry-wise non-negativity. We begin our analysis by considering the full-batch loss in Sec. 4.1, and extend our results to the SCL minimized on mini-batches in Sec. 4.2. We provide specific conditions that mini-batches must satisfy in order to have the same global optimum as the full-batch version.
@@ 4.1 Full-batch SCL
We defer the proof details to Sec. A in the appendix. The bound relies on successive uses of Jensen’s inequality and the fact that for feasible ${\mathbf{H}}$, each pair of training samples $i,j\in[n]$ satisfies $0\leq\mathbf{h}_{i}^{\top}\mathbf{h}_{j}\leq 1$. We complete the proof by verifying that equality is attained only if $\mathbf{h}_{i}^{\top}\mathbf{h}_{j}=1$ when $y_{i}=y_{j}$ and $\mathbf{h}_{i}^{\top}\mathbf{h}_{j}=0$ when $y_{i}\neq y_{j}$. Rather, to achieve the optimal cost, features with similar labels must align (NC) and the class-mean features must form an OF.
[…]
Figure 3: Full-batch SCL converges to the lower bound (dashed lines) in Thm. 1.
4.2 Mini-batch SCL
Thm. 1 shows that minimizing SCL over the training set as a single batch recovers an OF in the feature space, regardless of the training label distribution. However, in practice, SCL optimization is performed over batches chosen from training set [Kho+20, Gra+21]. Specifically, we have a set of batches $\mathcal{B}$ and we compute the loss on each batch $B\in\mathcal{B}$ as in (3). While in the full-batch SCL (4), all pairs of training samples interact with each other, in the mini-batch version, two samples $(i,j)$ interact only if there exists a batch $B\in\mathcal{B}$ such that $i,j\in B$. A natural question is whether the mini-batch construction impacts the embeddings learned. In this section, we study the role of batching on the embeddings geometry more closely.
where recall $n_{B,c}=|\{i:i\in B,y_{i}=c\}|$. Thm. 2 is an extension of Thm. 1 for the mini-batch SCL (10). The proof proceeds similarly to that of Thm. 1, where the loss over each batch can be independently bounded from below. Interestingly, due to the orthogonality of the optimal embeddings in every batch, the lower bound for each batch can be achieved by the common minimizer. The detailed proof is deferred to Sec. B in the appendix.
Theorem 2 (Mini-batch SCL minimizers).
[…]
Equality holds if and only if for every $B\in\mathcal{B}$ and every pair of samples $i,j\in B$, $\mathbf{h}_{i}^{\top}\mathbf{h}_{j}=0$ if $y_{i}\neq y_{j}$ and $\mathbf{h}_{i}=\mathbf{h}_{j}$ if $y_{i}=y_{j}$.
We remark that the only previous study of SCL geometry with batches by [Gra+21] requires significantly more restrictive conditions on the batch set ${\mathcal{B}}$ to guarantee a common geometry among all batches $B\in{\mathcal{B}}$. Specifically, [Gra+21] assumes a batch set that includes all possible combinations of examples. Instead, thanks to the addition of ReLU, our requirements are significantly relaxed. This is detailed in the following discussion.
@@ 4.2 Mini-batch SCL
Figure 4: Rem. 1 visualized, (a)(i,ii,iii) indicate the antipodal structure of the optimal embeddings under UFM in each of the 3 mini-batches respectively, whereas the overall optimal geometry is an ETF (a)(iv). This contrasts the optimal embeddings under ${}_{+}$ where each mini-batch (b)(i,ii,iii) is consistent with the overall optima (b)(iv).
Remark 1 (Comparison to analysis of [Gra+21, ]).
We elaborate on the comparison to [Gra+21, ] with an illustrative example below. When optimizing the loss decomposed over a set of mini-batches, it is crucial to consider whether the individual batch minimizers are also overall optimal solutions. Consider the following setting: $k=d=3,y_{1}=1,y_{2}=2,y_{3}=3,B_{1}=\{1,2\},B_{2}=\{1,3\},B_{3}=\{2,3\},\mathcal{B}=\{B_{1},B_{2},B_{3}\}$. Let us identify the optimal embeddings for the two scenarios: (i) with and (ii) without non-negativity constraints on the embedding coordinates.
(i) Without non-negativity (UFM), the optimal embedding configurations can be shown to be ETF with $2$ vectors for every batch, which is simply an antipodal structure. In other words, the batch-wise optimal solutions are $\mathbf{h}_{1}=-\mathbf{h}_{2},\mathbf{h}_{1}=-\mathbf{h}_{3}$ and $\mathbf{h}_{2}=-\mathbf{h}_{3}$, for the batches $B_{1},B_{2}$ and $B_{3}$, respectively. However, the batch-wise optimal solutions are infeasible due to their contradictory nature. From [Gra+21], it can be deduced that the overall optimal configuration is instead an ETF with $3$ vectors. This example underscores the difficulty in optimizing the loss over every batch separately in case of SCL without non-negativity. See (a) in Fig. 4.
(ii) With non-negativity ($\text{UFM}_{+}$), our results imply that the optimal embeddings form a $2$-OF for every batch, i.e., $\mathbf{h}_{1}\perp\mathbf{h}_{2},\mathbf{h}_{1}\perp\mathbf{h}_{3}$ and $\mathbf{h}_{2}\perp\mathbf{h}_{3}$, for the batches $B_{1},B_{2}$ and $B_{3}$, respectively. The three conditions are compatible with each other and the fact that the overall optimal solution is a $3$-OF in $\mathbb{R}^{3}$. Therefore, we were able to break down the overall optimization into individual batches. See (b) in Fig. 4.
Figure 5: A non-OF geometry minimizing mini-batch SCL with =$\mathcal{B}=\{B_{1},B_{2}\}$. Samples with different labels are marked with different colors.
Global minimizer geometry may not be unique. It is easy to verify that any ${\mathbf{H}}$ following the OF geometry achieves the lower-bound of Thm. 2 and it is also a global optimizer of the mini-batch SCL. However, depending on the choice of $\mathcal{B}$, the lower bound can possibly be attained by other optimal embeddings that do not necessarily satisfy NC or orthogonality. Hence the global optimizer may not have a unique geometry.
@@ 5 Batching matters
Since the OF may not be the only solution for the $\text{UFM}_{+}$ with mini-batch SCL (10), here we study the role of mini-batches to avoid ambiguous solutions with different geometries. In Sec. 5.1, we identify necessary and sufficient conditions for a batching strategy to yield a unique global solution geometry. By this result, in Sec. 5.2, we propose a simple yet effective scheme that provably succeeds in improving the convergence to an OF.
@@ 5.1 When is OF geometry the unique minimizer?
We state necessary and sufficient conditions on $\mathcal{B}$ for the minimizer of mini-batch SCL to be unique:
@@
Corollary 2.1 serves as a check for whether a given batching scheme yields a unique global minimizer geometry or not. It also provides guidance for designing mini-batches. We elaborate on this below.
5.2 Batch-binding ensures unique OF minimizer and improves convergence
Corollary 2.1 shows that an arbitrary set of mini-batches is not guaranteed to have OF geometry as its minimizing configuration in the embedding space. To enforce the OF geometry as the unique minimizer, we introduce a simple scheme with low computational overhead as shown in Alg. 1. See Fig. 15 in the appendix for a visual example.
@@ 5.2 Batch-binding ensures unique OF minimizer and improves convergence
Batch-binding adds a constant $k$ to the size of every mini-batch. $k$ is typically smaller than the batch sizes used to train SCL which could range from $2048$ to $6144$ [Kho+20]. While adding small computational overhead, this method guarantees that $\mathcal{B}^{\prime}$ satisfies the conditions of Cor. 2.1 and SCL learns an OF geometry. We conduct a series of experiments to illustrate the impact of batch binding on convergence to OF.
To verify the role of batch selection on the learned features, we train ResNet-18 on CIFAR-10 under three batching scenarios:
  1. 1.
    No batch-shuffling: a mini-batch set $\mathcal{B}$ that partitions the training examples once at initialization and is held constant across training epochs.
  2. 2.
    Batch-shuffling: the examples are divided into mini-batch partition, with a random reshuffle of the examples at every epoch.
  3. 3.
    No batch-shuffling + batch-binding: we construct a fixed partition of examples into mini-batches; then, we perform batch-binding. In order to focus on the impact of batch selection strategies, we consider a relatively small batch-size of $128$.
We remark that Thm. 2 and Cor. 2.1 can explain the behaviors of all three schemes. Moreover, the batch-binding strategy guided by our analytical results is effective at ensuring fast and predictable convergence of deep-nets to a unique OF geometry. Fig. 6 shows the distance of learned embeddings to the OF geometry for the three batching scenarios mentioned above. In case of the fixed partition batching (left), the features do not converge to OF, especially with large imbalance ratios. This behavior is consistent with the conditions identified by Thm. 2, because in a typical random partition, the corresponding induced subgraphs are not necessarily connected. On the other hand, when the batching is performed with a randomly reshuffled data ordering (center), the geometry converges more closely to the OF, albeit with certain epochs deviating significantly from the convergence direction. We hypothesize that random reshuffling creates an opportunity for different examples to interact sufficiently, thus, when trained long enough, to converge close to an OF geometry. Finally, with the third batching strategy of batch-binding (right), even without shuffling the samples at every epoch, we observe a consistently fast convergence to the global optimizer, OF. These observations provide strong evidence supporting our analysis of batching predicting the behavior of SCL trained deep-nets.
While the above studies the impact of batching itself, we perform a number of additional experiments in order to illustrate the impact of binding examples. In particular, we observe that batch binding helps improve convergence to OF geometry when training with less powerful models (see Fig. 18 for DenseNet experiments in the appendix), more complex dataset (see Fig. 19 in the appendix and Fig. 2 for CIFAR100 and TinyImageNet results) and even convergence to ETF in the absence of ReLU under balanced training (See Fig. 20 in the appendix). Such results emphasize the importance of batch binding and encourage further analysis of batching under contrastive learning.
@@
Figure 6: Convergence to the OF geometry for various batching schemes including the analysis-inspired scheme “No Batch Shuffling + Batch-Binding”. ResNet-18 trained on CIFAR-10 with a small batch size of 128, see text. See also Fig. 17 in the appendix
6 More Related Work